OpenStudy (annie96):

Graph the solution of the open sentence. –5 ≤ n – 2 ≤ 2.

OpenStudy (anonymous):

well what do you get when you leave n by itself

OpenStudy (anonymous):

youll get \[-3\le n \le 4\]

OpenStudy (anonymous):

so if you seperate it youll get two line which is \[-3\le n \]\[n \le 4\]

OpenStudy (anonymous):

you can just plug it in the chart and youll get the answer

OpenStudy (anonymous):

and also because the signs are greater than or equal and less than or equal the dot has to be closed

OpenStudy (annie96):

k thank you

OpenStudy (anonymous):

so the answer will be C